# Weighted algebraic topology, II

### 05 May 2026 v1

@@ 6.3. Corollary I (Reflective symmetrisation and products)
(a) For a finite family of $\mu$-metric spaces $X_{1},...,X_{n}$:
@@
(b) More particularly, for $\delta$-spaces:
@@ 6.4. Corollary II (Coreflective symmetrisation and products)
(a) For a finite family of $\mu$-spaces $X_{1},...,X_{n}$:
@@
Therefore the functor $^{\,{\scriptstyle{\vee}}\,}\colon\mu_{\infty}\mathsf{Mtr}\to\mathsf{Top}$ preserves finite products (but takes arbitrary products to the box topology).
(b) More particularly, for $\delta$-spaces:
@@
(c) If the symmetrisations $X_{i}^{\,{\scriptstyle{\vee}}\,}$ of the $\mu$-spaces $X_{i}$ are positive:
@@
and again these $\mu$-spaces are Lipschitz-equivalent (but $(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}$ need not be).
@@
(b) As in Corollary I, the inequality $(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}\leqslant(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}$ follows from $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}\leqslant\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i}$, proved in (83) for $\delta$-spaces. We know that this does not extend to $\mu$-spaces, even in the symmetric case.
@@ 6.5. Ascent and descent
(a) Re-examining the elementary example of the Introduction, in 0.1, we consider the $\mu$-space $X=\delta\mathbb{R}\!\times\!\mu\mathbb{R}$, with the product $\mu$-metric $\mu((x,y),(x^{\prime},y^{\prime}))=\delta(x,x^{\prime}){\,{\scriptstyle{\vee}}\,}\mu(y,y^{\prime})$
(96) $\mu((x,y),(x^{\prime},y^{\prime}))\;=\;\begin{cases}(x^{\prime}-x){\,{\scriptstyle{\vee}}\,}(y^{\prime}-y),&\;\mbox{ if }\>x\leqslant x^{\prime},\\[3.0pt] \infty&otherwise.\end{cases}$
(b) A $\mu_{\infty}$-map $a=\langle a_{1},a_{2}\rangle\colon\delta\mathbb{I}\to X$
@@
has components $a_{1},a_{2}$ such that:
@@
- the path $a_{2}\colon\delta\mathbb{I}\to\mu\mathbb{R}$ satisfies the Lipschitz condition, examined in 4.6 and 5.1(d).
(c) Evaluation, total ascent, total descent and total variation have already been examined in 0.1, consistently with the definitions of 5.2.
@@ 6.6. Gravitational potential and metric
We rewrite it in the form:
(99) $\varphi(x)=k-k/(1+x)=kx/(1+x),\qquad\;\;\mbox{for }x\geqslant 0,$
where $x=r/r_{0}-1\geqslant 0$ is the altitude, measured with respect to the Earth radius; we have added the constant $k$ so that the potential $\varphi$ annihilates at the Earth surface: $\varphi(0)=0$. The function $\varphi$ is increasing.
The associated (linear) $\mu$-metric $\mu(x,y)=\varphi(y)-\varphi(x)$ measures the relative gain of potential energy from altitude $x$ to altitude $y$, or the opposite of the work of the gravitational field in this transition.
(b) At small altitude, that is, for small $x$, one can use a linear approximation $\psi(x)=kx$ (which amounts to taking the scalar gravitational field constant), getting a $\mu$-metric
(100) $\mu^{\prime}(x,y)=\psi(y)-\psi(x)=k(y-x)\qquad\;\;\mbox{for }x,y\geqslant 0,$
@@ 6.7. Elastic potential and metric
(102) $\varphi(x,y)=\lambda(x^{2}+y^{2}),\qquad\mbox{for }(x,y)\in\mathbb{R}^{2},$
@@
where $k=2\lambda>0$ is the elastic constant of the field: F(x,y)=k(x,y)F(x,y)=-k(x,y). The associated (linear) $\mu$-metric
(103) $\mu((x,y),(x^{\prime},y^{\prime}))=\varphi(x^{\prime},y^{\prime})-\varphi(x,y)=\lambda((x^{\prime 2}+y^{\prime 2})-(x^{2}+y^{2})),$
@@
measures the relative gain of potential energy from position $(x,y)$ to $(x^{\prime},y^{\prime})$, that is the opposite of the work of the elastic field in this transition.
@@
Letting $\mu_{E}\mathbb{R}$ have $\mu_{E}(x,x^{\prime})=x^{\prime 2}-x^{2}$, we are considering the linear $\mu$-space $(\lambda\mu_{E}\mathbb{R})^{\otimes 2}$.
7. Relative antimetrics and Minkowski spacetime
Reversing inequalities, we briefly consider antimetrics, already mentioned in 2.1(b). The theory is obviously equivalent to the previous one, but more natural in special relativity. The future timecone of a Lorentz vector space has a well known ‘antinorm’, which we use in 7.4 to define an antimetric $\gamma(x,y)$ of the whole spacetime; the latter is $\geqslant 0$ if the event $x$ can influence $y$, and $\overline{\infty}$} otherwise. The associated relative metric is $\mu(x,y)=-\gamma(x,y)$. Geodesic paths have thus a maximal valuation with respect to $\gamma$ (the longest elapsed proper time), and minimal (if negative) with respect to the relative metric.
We essentially follow the terminology of O’Neill’s book on semi-Riemannian manifolds [On], although here it would be more natural to use the opposite signature of the metric tensor.
7.1. Relative antimetrics
(a) An antimetric on a set $X$ is a mapping $\gamma\colon X\!\times\!X\to\overline{\mathbb{R}}$ satisfying the axioms
(104) $\gamma(x,y)+\gamma(y,z)\leqslant\gamma(x,z),\qquad 0\leqslant\gamma(x,x)\qquad\;\;(x,y,z\in X).$
The relative value $\gamma(x,y)$ represents the gain, or profit, of the transition from $x$ to $y$ – which is unaffordable, or forbidden, when $\gamma(x,y)={\overline{\infty}}$. The reverse triangular inequality expresses the fact that a combined transition should grant economies of scale; $\gamma(x,x)$ can only be 0 or $\infty$. Plainly, a linear metric is also a (linear) antimetric.
Formally, $(X,\gamma)$ is a small category enriched over the symmetric monoidal closed category $\bf w$}* associated to the quantale $\overline{\mathbb{R}}$, with its natural order and the extended sum where $\infty+{\overline{\infty}}={\overline{\infty}}$ (so that the initial object $\overline{\infty}$} is preserved by adding $\infty$).
(b) The object $X=(X,\gamma)$ will also be called a $\gamma$-space. The previous theory of $\mu$-spaces can be simply rewritten for these structures, reversing inequalities, $\rm inf$} and $\rm sup$}, $\infty$ and $\overline{\infty}$}.
A Lipschitz map $f\colon X\to Y$ of $\gamma$-spaces has a Lipschitz constant $\lambda\in[0,\infty$[ such that
(105) $\lambda\,\gamma_{X}(x,x^{\prime})\,\leqslant\,\gamma_{Y}(f(x),f(x^{\prime})),\qquad\mbox{for }x,x^{\prime}\in X.$
They form the category $\gamma_{\infty}\mathsf{Mtr}$ of $\gamma$-spaces and Lipschitz maps. For $\lambda=1$ we have the wide subcategory $\gamma\mathsf{Mtr}$ of gain-augmenting maps.
From 5.2, a set-theoretical path $a\colon\mathbb{I}\to X$ has a gain-valuation $v_{\gamma}(a)$ defined as follows (with the usual meaning of a sequence ${\underline{t}}:\;0=t_{0}\leqslant t_{1}\leqslant...\leqslant t_{p}=1$)
(106) $v_{\underline{t}}(a)=\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\gamma(a(t_{j-1}),a(t_{j})),\qquad v_{\gamma}(a)={\rm inf}_{\underline{t}}\,v_{\underline{t}}(a)\in\overline{\mathbb{R}}.$
(c) Let $X$ be an abelian group. Applying 3.3, an antimetric $\gamma$ on $X$ is invariant up to translations if and only if it can be expressed as $\gamma(x,y)={\underline{\gamma}}(y-x)$, by an antinorm ${\underline{\gamma}}\colon X\to\overline{\mathbb{R}}$, satisfying the axioms
(107) ${\underline{\gamma}}(x)+{\underline{\gamma}}(y)\,\leqslant\,{\underline{\gamma}}(x+y),\qquad 0\leqslant{\underline{\gamma}}(0)\qquad\;\;(x,y\in X),$
Note that ${\underline{\gamma}}(0)$ can only be 0 or $\infty$.
7.2. The Lorentz vector space
(a) The n-dimensional Lorentz vector space $V=\mathbb{R}^{n}_{1}$ is a real vector space of dimension $n\geqslant 2$, equipped with the Lorentz scalar product $\bf g$}, of index 1
(108) ${\bf g}(v,w)=\langle v,w\rangle=\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{i}\,\varepsilon_{i}v_{i}w_{i}\qquad\;\;(i=1,...,n),$
and signature $\underline{\varepsilon}=(-1,1,...,1)$. $V$ can be viewed as the tangent vector space of the Minkowski spacetime, at any point.
The scalar product is a non-degenerate symmetric bilinear mapping $V\!\times\!V\to\mathbb{R}$. We recall that ‘non-degenerate’ means that $\langle v,w\rangle$ only annihilates for all $w$ if $v=0$ (as shown letting $w$ vary in the canonical base $e_{1},...,e_{n}$ of $\mathbb{R}^{n}$).
The associated quadratic form
(109) ${\bf q}(v)=\langle v,v\rangle=\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{i}\,\varepsilon_{i}v_{i}^{2},$
is indefinite; it will be used below to define an antinorm on $V$.
(b) For a vector $v\in V$, the number $v_{1}$ is the time component of $v$, while the vector ${\vec{v}}=(v_{2},...,v_{n})\in\mathbb{R}^{n-1}$ is its spatial component; we write as ${\vec{v}}||$ the euclidean norm of the latter
[…]
7.3. An antinorm for the Lorentz vector space
(a) The closed future cone of $V$ is the following subset (related to the future causal cones of the spacetime, in 7.4)
[…]
Note that $\lambda v+\mu w\in\Lambda$, for all $v,w\in\Lambda$ and $\lambda,\mu\geqslant 0.$ If $v,w\in\Lambda$ then $\langle v,w\rangle\leqslant 0$, as it follows applying the (ordinary) Cauchy-Schwarz inequality to the spatial part
[…]
(b) The open future cone of $V$ is its interior $\Lambda^{\circ}$, for the euclidean topology (and corresponds to the future timecones of 7.4.)
[…]
$\Lambda^{\circ}$ is the timecone $C(e_{1})$, in the notation of [On], p. 143. The set of timelike vectors, defined by the inequality $\langle v,v\rangle<0$, is the disjoint union of $\Lambda^{\circ}$ with the opposite timecone $\Lambda^{\circ}=C(-e_{1})$.
[…]
(d) We define now an antinorm on the vector space $V$ letting
[…]
(e) We have to prove the reverse triangular inequality for $\underline{\gamma}$}. On the cone $\Lambda^{\circ}$ this is an important, well known fact (see [On], Corollary 5.31), related to the ‘reverse Schwarz inequality’. Its extension to $V$ (once the value $\overline{\infty}$} is allowed) is easy.
First, the inequality holds on the closed cone $\Lambda$: if $v$ or $w$ is in $\partial\Lambda$, taking into account that $\langle v,w\rangle\leqslant 0$ and ${\underline{\gamma}}(v)\,{\underline{\gamma}}(w)=0$
$\begin{array}[]{c}({\underline{\gamma}}(v+w))^{2}=-\langle v+w,v+w\rangle=-\langle v,v\rangle-2\langle v,w\rangle-\langle w,w\rangle\\[5.0pt] \geqslant\,{\underline{\gamma}}(v)^{2}+{\underline{\gamma}}(w)^{2}=({\underline{\gamma}}(v)+{\underline{\gamma}}(w))^{2}.\end{array}$
Finally, if one at least of $v,w$ does not belong to $\Lambda$, we trivially have ${\underline{\gamma}}(v)+{\underline{\gamma}}(w)={\overline{\infty}}\leqslant{\underline{\gamma}}(v+w)$.
7.4. An antimetric for the spacetime
(a) A Minkowski spacetime $M$ is a time-oriented $n$-dimensional Lorentz manifold that is isometric to the Minkowski n-space $X=\mathbb{R}^{n}_{1}$, for $n\geqslant 2$ (slightly adapted from [On], Chapter 6, Definition 8).
We shall identify $M=X$, adopting an inertial coordinate system in $M$ that preserves time-orientation ([On], Chapter 6, Definition 11). At any $x\in X$, the tangent vector space $T_{x}X$ can be identified with the Lorentz vector space $V=\mathbb{R}^{n}_{1}$ examined above, with (indefinite) scalar product ${\bf g}(v,w)=\langle v,w\rangle=\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{i}\,\varepsilon_{i}v_{i}w_{i}$.
$V$ and $X$ have thus the same underlying set, but a different structure:
- $V$ is a Lorentz vector space, and we still write its vectors as v,w,,v,w,...,
- $X$ is a semi-Riemannian manifold, with Lorentz vector space $T_{x}X=V$ at any point; its elements, written as x,y,,x,y,..., are called events.
$V$ acts on the affine space $X$ by the sum $x+v$ (of $V$). A difference yxy-x of events is interpreted as the vector $v\in V=T_{x}X$ such that $x+v=y$.
(b) Given two distinct events $x,y$, we recall that $y$ can be reached by a material particle (resp. a lightlike particle) starting at $x$ if and only if it belongs to the future timecone $x+\Lambda^{\circ}$ (resp. the future lightcone $x+\partial\Lambda$, excluding $x$). Globally, the event $y$ can be influenced by $x$ if and only if it belongs to the future causal cone $C^{+}(x)=x+\Lambda$.
(c) The antinorm ${\underline{\gamma}}\colon V\to\overline{\mathbb{R}}$ defined in 7.3(d) gives an antimetric
(114) $\gamma\colon X\!\times\!X\to\overline{\mathbb{R}},\qquad\qquad\gamma(x,y)={\underline{\gamma}}(y-x).$
Keeping $x$ fixed, $\gamma(x,y)$ is finite and (weakly) positive when $y$ belongs to the future causal cone $C^{+}(x)=x+\Lambda$, but takes value $\overline{\infty}$} on all events $y$ which cannot be influenced by $x$. It annihilates on the future lightcone $x+\partial\Lambda$.
7.5. Lipschitz paths in the spacetime
The relative metric of the Minkowski spacetime $M=X$ is thus
(115) $\mu(x,y)=-\gamma(x,y),$
and takes values in ${\overline{\infty}},0]\cup\{\infty\}$. (Let us note that the gain function $\gamma$, which is positive where it matters, is more consistent with the usual interpretation of spacetime.)
The reflective (pre)order in the $\mu$-space $M$ is the causality order, defined by the future causal cones:
(116) $x\leqslant y\;\iff\;\mu(x,y)<\infty\;\iff\;y\in C^{+}(x)=x+\Lambda.$
It is easy to see that a mapping $a\colon\delta\mathbb{I}\to M$ is a Lipschitz path if and only if it is weakly increasing for the reflective orders of these $\mu$-spaces:
(117) $t$ $\leqslant t^{\prime}\mbox{ in }\mathbb{I}\;\implies\;a(t)\leqslant a(t^{\prime})\mbox{ in }M$.
Special relativity is interested in the piecewise differentiable paths $a\colon\mathbb{I}\to M$ that satisfy this condition, and can thus represent the movement of a material or lightlike particle. Moreover, the integral of the metric tensor along the path measures the proper duration of the path itself – formally zero for a light ray (which can bear no clock).
Remarks. (a) The reflective symmetric $\mu$-metric of $M$ is chaotic: $\hat{\mu}(x,y)={\overline{\infty}}$ for all $x,y$. In fact, we can take an event $z=x+\lambda e_{1}$ with $\lambda>0$ sufficiently big so that $z\in C^{+}(y)$.
Then $\mu(x,y)=-{\underline{\gamma}}(\lambda e_{1})=-\lambda\;$ and $\mu(y,z)\leqslant 0$, whence
$\hat{\mu}(x,y)\leqslant\hat{\mu}(x,z)+\hat{\mu}(z,y)\leqslant\mu(x,z)+\mu(y,z)\leqslant-\lambda,$
and this will still hold for any $\lambda^{\prime}\geqslant\lambda$.
(b) The coreflective symmetric $\mu$-metric of $M$ is obviously discrete: given two distinct events, one of them at least cannot be influenced by the other.
7.6. Back to topologies
Finally, all this supports our claim: there is no sensible way to derive a definite topology from a relative metric.
The Minkowski spacetime $M$ has a natural topology, defined by the manifold structure and homeomorphic to the euclidean space $\mathbb{R}^{n}$. The ‘acceptable’ paths $a$ in $M$, described above, are determined by the differentiable structure and the metric tensor (or the deduced $\mu$-metric); they are continuous for the natural topology.
On the other hand, the reflective topology associated to $\mu$ is chaotic, and the coreflective one is discrete, as we have already seen.
The future topology associated to $\mu$ is of some (slight) interest: the $\varepsilon$-ball of an event $x$ always coincides with the whole causal cone $C^{+}(x)$, which is thus determined by ${\rm tp}^{+}(M)$; yet, continuous paths for this topology are quite erratic, only respecting condition (117).

### 13 May 2026 v2

@@ 6.3. Corollary I (Reflective symmetrisation and products)
(a) For a finite family of $\rho$-metric spaces $X_{1},...,X_{n}$
@@
(b) More particularly, for $\delta$-spaces
@@ 6.4. Corollary II (Coreflective symmetrisation and products)
(a) For a finite family of $\rho$-spaces $X_{1},...,X_{n}$
@@
Therefore the functor $^{\,{\scriptstyle{\vee}}\,}\colon\rho_{\infty}\mathsf{Mtr}\to\mathsf{Top}$ preserves finite products (but takes arbitrary products to the box topology).
(b) More particularly, for $\delta$-spaces
@@
(c) If the symmetrisations $X_{i}^{\,{\scriptstyle{\vee}}\,}$ of the $\rho$-spaces $X_{i}$ are positive
@@
and again these $\rho$-spaces are Lipschitz-equivalent (but $(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}$ need not be).
@@
(b) As in Corollary I, the inequality $(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}\leqslant(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}$ follows from $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}\leqslant\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i}$, proved in (76) for $\delta$-spaces. We know that this does not extend to $\rho$-spaces, even in the symmetric case.
@@ 6.5. Ascent and descent
(a) Re-examining the elementary example of the Introduction, in 0.2, we consider the $\rho$-space $X=\delta\mathbb{R}\!\times\!\rho\mathbb{R}$, with the product $\rho$-metric $\rho((x,y),(x^{\prime},y^{\prime}))=\delta(x,x^{\prime}){\,{\scriptstyle{\vee}}\,}\rho(y,y^{\prime})$
(89) $\rho((x,y),(x^{\prime},y^{\prime}))\;=\;\begin{cases}(x^{\prime}-x){\,{\scriptstyle{\vee}}\,}(y^{\prime}-y),&\;\mbox{ if }\>x\leqslant x^{\prime},\\[3.0pt] \infty&otherwise.\end{cases}$
(b) A $\rho_{\infty}$-map $a=\langle a_{1},a_{2}\rangle\colon\delta\mathbb{I}\to X$
@@
has components $a_{1},a_{2}$ such that
@@
- the path $a_{2}\colon\delta\mathbb{I}\to\rho\mathbb{R}$ satisfies the Lipschitz condition, examined in 4.6 and 5.1(d).
(c) Evaluation, total ascent, total descent and total variation have already been examined in 0.2, consistently with the definitions of 5.2.
@@ 6.6. Gravitational potential and metric
We rewrite it in the form
(92) $\Phi(x)=k-k/(1+x)=kx/(1+x),\qquad\;\;\mbox{for }x\geqslant 0,$
where $x=r/r_{0}-1\geqslant 0$ is the altitude, measured with respect to the Earth radius; we have added the constant $k$ so that the potential $\Phi$ annihilates at the Earth surface: $\Phi(0)=0$. The function $\Phi$ is increasing.
The associated (linear) $\rho$-metric $\rho(x,y)=\Phi(y)-\Phi(x)$ measures the gain of potential energy from altitude $x$ to altitude $y$, or the opposite of the work of the gravitational field in this transition.
(b) At small altitude, that is, for small $x$, one can use a linear approximation $\psi(x)=kx$ (which amounts to taking the scalar gravitational field constant), getting a $\rho$-metric
(93) $\rho^{\prime}(x,y)=\psi(y)-\psi(x)=k(y-x)\qquad\;\;\mbox{for }x,y\geqslant 0,$
@@ 6.7. Elastic potential and metric
(95) $\Phi(x,y)=\lambda(x^{2}+y^{2}),\qquad\mbox{for }(x,y)\in\mathbb{R}^{2},$
@@
where $k=2\lambda>0$ is the elastic constant of the field: F(x,y)=k(x,y)F(x,y)=-k(x,y). The associated (linear) $\rho$-metric
(96) $\rho((x,y),(x^{\prime},y^{\prime}))=\Phi(x^{\prime},y^{\prime})-\Phi(x,y)=\lambda((x^{\prime 2}+y^{\prime 2})-(x^{2}+y^{2})),$
@@
measures the gain of potential energy from position $(x,y)$ to $(x^{\prime},y^{\prime})$, that is the opposite of the work of the elastic field in this transition.
@@
Letting $\rho_{E}\mathbb{R}$ have $\rho_{E}(x,x^{\prime})=x^{\prime 2}-x^{2}$, we are considering the linear $\rho$-space $(\lambda\rho_{E}\mathbb{R})^{\otimes 2}$.